3. An analyst gathers or estimates the following information about a stock:
- Current price per share: $15.5
- Current annual dividend per share: $1.6
- Annual dividend growth for Years 1-3: 8%
- Annual dividend growth for Years 4+: 3%
- Required rate of return: 11%
Based on the DDM, the stock is most likely:
A. undervalued. B. overvalued. C. fairly valued.

Please Don't solve it in Excel.
Answer #1

Answer

Stick price using DDM dividend discount model.

Current dividend( D​​​​​0)=1.6

Annual dividend growth upto 3 years=8%

Therefore

D1( dividend after 1 year)=1.6(1+0.08)=1.728

D2( dividend after 2 year)=1.728(1+0.08)=1.866

D3(dividend after 3 years)=1.866(1+0.08)=2.015

P3( price after 3 years)=

Price = D3(1+g)/(Re- g)=2.015(1.04)/(11-4)=29.9 approx

Re= Required rate of return=11%

growth after 3 years=4%

Therfeore price of stock is 29.9 as per dividend discount model.

current price per share=15.5

Based on DDM stock is undervalued.